Create a table to represent your table visual basics Module.

  1. Open a blank Word or Text file, save this file as: yourlastname-ProgrammingLooping.docx or yourlastname-ProgrammingLooping.txt. Then type the answer to the following questions:
  2. What is the output of the following codes, consider each case independently. I have an example of Case 1 to show you how to complete each case at the end of this assignment.

This is the Question to be anwsered Case 1: What will be the value of x after the following loop is executed:

x = 0

For i = 1 To 2

For j = 1 To 3

     x = x * i * j

Next j

Next i

NOTE:

This is an example of what your anwser should look like:

x

i

j

NOTES

0

x is initialized to 0

x = 0*1*1, the new value for x is 0

1

1

First we start with the outer loop (for I ) and continue with inner loop (for J ).

x = 0*1*2, the new value for x is 0

1

2

Still in the inner loop when J becomes 2, note I is 1

x = 0*1*3, the new value for x is 0

1

3

Still in the inner loop when J becomes 3, note I is 1

x = 0*2*1, the new value for x is 0

2

1

Here program continues with the outer loop, where I is 2 and moves to inner loop where J is 1

x = 0*2*2, the new value for x is 0

2

2

Program continues with the inner loop when J becomes 2, note I is 2

x = 0*2*3, the new value for x is 0

2

3

Program continues with the inner loop when J becomes 3, note I is 2

X = 0

Is the final value for x when both loops are completed
Answer #1

Answer :

x

i

j

NOTES

0

x is initialized to 0

x = 0*1*1, the new value for x is 0

1

1

First we start with the outer loop (for I ) and continue with inner loop (for J ).

x = 0*1*2, the new value for x is 0

1

2

Still in the inner loop when J becomes 2, note I is 1

x = 0*1*3, the new value for x is 0

1

3

Still in the inner loop when J becomes 3, note I is 1

x = 0*2*1, the new value for x is 0

2

1

Here program continues with the outer loop, where I is 2 and moves to inner loop where J is 1

x = 0*2*2, the new value for x is 0

2

2

Program continues with the inner loop when J becomes 2, note I is 2

x = 0*2*3, the new value for x is 0

2

3

Program continues with the inner loop when J becomes 3, note I is 2

X = 0

Is the final value for x when both loops are completed